### v1

# Diffusion in Zero-Shot Learning for Environmental Audio

Abstract
Zero-shot learning enables models to generalize to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes. While much research has focused on zero-shot learning in computer vision, the application of these methods to environmental audio remains underexplored, with poor performance in existing studies. Generative methods, which have demonstrated success in computer vision, are notably absent from environmental audio zero-shot learning, where classification-based approaches dominate.
To address this gap, this work investigates generative methods for zero-shot learning in environmental audio. Two successful generative models from computer vision are adapted: a cross-aligned and distribution-aligned variational autoencoder (CADA-VAE) and a leveraging invariant side generative adversarial network (LisGAN). Additionally, a novel diffusion model conditioned on class auxiliary data is introduced. The diffusion model generates synthetic data for unseen classes, which is combined with seen-class data to train a classifier.
Experiments are conducted on two environmental audio datasets, ESC-50 and FSC22. Results show that the diffusion model significantly outperforms all baseline methods, achieving more than 25% higher accuracy on the ESC-50 test partition.
This work establishes the diffusion model as a promising generative approach for zero-shot learning and introduces the first benchmark of generative methods for environmental audio zero-shot learning, providing a foundation for future research in the field.
Code is provided at https://github.com/ysims/ZeroDiffusion for the novel ZeroDiffusion method.
Index Terms:
Environmental Audio, Zero-Shot Learning, Generative Networks
I Introduction
Machine learning has revolutionised how society solves complex problems [1, 2, 3]. Various publications consider the future impact of machine learning, such as AI cities [4], humanoid robotics [5] and medicine [6, 7, 8]. To solve these future problems, machines and AI need to be adaptive [9, 10]. One way machines can be adaptive is to classify new items without labelled training data [11, 12]. This problem is called ‘zero-shot learning’.
Zero-shot learning is a category of supervised machine learning that aims to create a model to classify classes it has not seen during training. That is, the set of classes used in training and testing are disjoint. Zero-shot learning is valuable in real-world applications, such as remote sensing [13], image classification [14, 15], image segmentation [16, 17], environmental audio [18, 19] and security [20], where there may not exist data for every class.
An intuitive way to understand zero-shot learning involves a thought experiment on how humans would solve the problem. Suppose someone has never seen a snake, cat or bird but knows key attributes these animals have. When shown a picture of a cat, the person reasons that it is a cat as it has four legs and whiskers, and no scales or wings. The person performs zero-shot learning using features of the target class and learning what those features mean through knowledge of other animals.
It is important not to confuse zero-shot learning with similar zero-data problems. In zero-shot learning, the method knows about the existence of all classes, i.e. the set of classes is closed. If it was open, i.e. there are unknown classes, the problem is called open-set learning [21, 22, 23]. Domain adaptation [24, 25] is another problem with a lack of data; however, there is no clear ‘seen’ and ‘unseen’ class distinction. An example of domain adaptation is training in simulation and transferring to the real world without fine-tuning [26, 27, 28]. Finally, zero-shot prompting is a newer area that uses large language models to perform new tasks [29, 30, 31]. This differs from zero-shot learning as there is no seen and unseen class distinction, and it is hard to quantify how much overlap the task has to scenarios the model was trained on.
Close to these zero-data problems, particularly zero-shot learning, is few-shot learning. Rather than zero data for select classes, few-shot learning has a small number of samples (e.g. three to five) for particular target classes. A comprehensive review of few-shot learning is provided by Song et al. [32].
In the literature, zero-shot learning methods are commonly tested on computer vision classification tasks [12, 33, 34, 14, 35, 36, 15, 37, 38, 39]. Traditional methods for zero-shot learning in computer vision relied on classification-based models with engineered attributes [12], but more recent methods use generative models [34, 15, 37] such as generative adversarial networks [40] and variational autoencoders [41]. These models have been shown to improve performance on unseen classes and can help mitigate problems such as the hubness problem and the domain shift problem [42]. The hubness problem occurs when data points are concentrated around a minority of classes, leading to poor performance when tested on unseen classes [43]. The domain shift problem arises when the model performs poorly because the training and test sets have different distributions [44].
This work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision. It takes inspiration from the successful usage of generative methods in computer vision [14, 45, 34, 37], and from a newer generative method, diffusion, which has seen success in image generation tasks [46, 47]. There is currently no literature known to the authors on environmental audio zero-shot learning using generative methods.
Recent use cases for environmental audio classification include smart cities [48], detection of traffic noise pollution [49], and hearing aid analysis [50]. These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
There is a lack of datasets that can be used to benchmark environmental audio zero-shot learning, as datasets are often not released in their raw audio form or lack classes. AudioSet [51] is a valuable large dataset, but instead of audio clips, it contains embeddings. Embeddings are high dimensional vectors representing the data. These embeddings come from a network pre-trained on AudioSet. This pre-training makes the dataset unsuitable for zero-shot learning, as the unseen classes may be present. Other datasets, such as UrbanSound8k [52], have too few classes for rigorous zero-shot learning. Partitioning seen and unseen classes will result in a seen set with little diversity and may not cover a large enough semantic space for the model to generalise to unseen classes, risking overfitting. Additionally, the unseen classes will be very few, and so the statistical significance of the results is poor. A reasonable minimum may be five unseen classes, and double that amount of seen classes, for a total of 15 classes in the dataset.
ESC-50 [53] is an open dataset with fifty classes, often used in environmental audio zero-shot learning literature. FSC22 [54] is another open dataset and contains twenty-seven classes. While this is almost half the size of ESC-50, it is still a substantial size to perform zero-shot learning on. We are not aware of the existence of any zero-shot learning literature using this dataset.
The present study uses ESC-50 and FSC22 to compare three generative methods. These methods use the following architectures: generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion networks. A bilinear compatibility network [55] is also compared as a baseline. The method ZeroDiffusion is introduced as a new method for zero-shot learning using diffusion.
The GAN method used is LisGAN [15]; however, the component where high-confidence unseen samples are injected back into the training is not included, to prevent violating the zero-shot learning problem formation where unseen class data should not be present during training. The method uses a conditional Wasserstein GAN [56]. Zero-shot learning methods often aim to learn the relationship between feature and class embeddings. To achieve this, the GAN is conditioned on class embeddings. It is then used to generate synthetic samples for unseen classes. Vanilla conditional GAN methods were experimented with in the presented study, but they suffered extreme instability and low performance. The adjusted LisGAN method performs worse than the bilinear compatibility method but is better than a random model.
The cross-aligned and distribution-aligned VAE (CADA-VAE) [14] has been previously used in computer vision with success. The method trains two variational autoencoders, one for auxiliary data and one for feature data, to learn a shared latent space. The results were comparable to the bilinear compatibility method.
Finally, the most successful model is our new diffusion model. The experiments with diffusion show a clear improvement upon any previous method in environmental audio zero-shot learning. While there is existing zero-shot learning literature outside of environmental audio on diffusion models [57, 58], no method trains a network from scratch. The methods use the pre-trained Stable Diffusion model [47]. While this can provide practical usage, it is hard to draw theoretical conclusions on the zero-shot performance of this method. The overlap of the unseen classes and training classes is unknown. This work shows that diffusion, with a transparent training process and a valid seen and unseen class distinction, outperforms other techniques in zero-shot learning.
The contributions of the presented paper are:
  • The first diffusion method for zero-shot learning that does not use a pre-trained model.
  • A new state-of-the-art for zero-shot learning with the ESC-50 dataset using the diffusion method.
  • The first benchmark on the FSC22 dataset, with the diffusion method outperforming other methods.
  • The first comparison of GANs, VAEs and diffusion-based methods for environmental audio zero-shot learning with ESC-50 and FSC22.
II Related Work
This section first covers terminology in zero-shot learning to better understand the literature. Then, it examines existing work on generative methods in zero-shot learning, much of which is in the domain of computer vision. Finally, existing works in environmental audio zero-shot learning are considered.
II-A Terminology
In this section, zero-shot learning terminology is presented in an effort to contextualise the literature. First, embeddings for both class and feature data are considered, then the concept of auxiliary data and finally, a high-level definition of zero-shot learning.
II-A1 Embeddings
Embeddings are a method of converting raw data, such as images or audio, into a high-level form that machine learning algorithms can use. The form embeddings take is a high-dimensional vector of real numbers. The methods in this work use deep learning to obtain these vector representations, using both supervised and unsupervised methods. An example of what an embedding space might look like is shown in Figure 1. Lions and cats are both felids and therefore semantically similar and closer to each other in the vector space. ‘Dog’ is semantically closer to the felids than ‘car’ as they are types of animals. In this way, the structure of the space reveals semantic relationships.
Figure 1: An embedding space. Semantically similar data is closer, such as ‘lion’ and ‘cat’.
Supervised methods to create embeddings typically involve classical supervised learning on a dataset and the output of the second last layer of the artificial neural network becomes the embedded data vector. Zero-shot learning requires a clear distinction between the seen and unseen classes, and so the dataset used to learn the embedding must only contain classes that are not part of the unseen class set. Typically the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset could be used as long as there is no overlap with the unseen class set.
Unsupervised methods involve learning a mapping to a space where the semantic distance of data is represented in the space. The most common example of this embedding method is word embeddings, such as GloVe [59] and Word2Vec [60]. These methods learn the semantic similarity of words using their proximity in texts. To a human, embeddings may seem to contain little meaning, but these vectors are information-rich. For example, the word embedding vectors for ‘man’, ‘woman’ and ‘king’ can be combined in an equation ‘king’ - ‘man’ + ‘woman’ to approximate the vector for ‘queen’ [60].
II-A2 Auxiliary Data
Zero-shot learning involves learning to classify known classes without samples for those classes. If the classes are unknown, the problem is called open-set learning. These known classes contain information even without samples. This information is called ‘auxiliary data’. The most common form of auxiliary data is the word embedding of the class label, retrieved using Word2Vec or GloVe. Another popular form is human-engineered attributes, where a vector is created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’. The Animals with Attributes [12] and Animals with Attributes 2 [61] datasets are examples of datasets that contain human-engineered attribute vectors, and were created for zero-shot learning research.
II-A3 Zero-Shot Learning
Zero-shot learning involves training on seen classes and testing on unseen classes. Generalised zero-shot learning extends this problem to include both seen and unseen classes in testing. Zero-shot learning is one of many zero-data problems, as described in the introduction.
Feature data is defined as the target data the user wants to classify. In computer vision, the feature data is images. In environmental audio, it is audio clips. Feature data for unseen classes should not be present anywhere before the testing phase, including during pre-training.
The embeddings for the feature data and classes represent semantic information. Zero-shot learning aims to train a model that generalises its understanding of semantics to go beyond class-specific knowledge.
II-B Generative Methods in Zero-Shot Learning
While there is no literature known to the authors in environmental audio that considers generative networks, there are many examples in the domain of computer vision. These largely involve networks combining GANs and autoencoders.
In 2017, autoencoders entered the computer vision zero-shot learning space through two publications, Tsai et al. [33] and Mukherjee et al. [62]. These used two traditional autoencoders for each modality and learnt a shared latent space between these two autoencoders. Conditional variational autoencoders were introduced in zero-shot learning the following year [34]. A new benchmark was found when CADA-VAE [14] was proposed which took advantage of variational autoencoders and created a shared latent space through cross-alignment and distribution-alignment.
One of the first GAN methods in zero-shot learning is f-CLSWGAN, a Wasserstein GAN method [36]. In the same year, another GAN method was created that used a visual pivot regulariser to guide training [63]. It also used Wikipedia text descriptions converted to word embeddings using term frequency-inverse document frequency (TF-IDF). The method was compared with other zero-shot learning methods including a basic GAN model. The basic GAN model performs much worse than any established zero-shot learning method, while Zhu et al.’s method [63] shows an improvement over the state-of-the-art.
LisGAN [15] improved on previous GAN methods by using a Wasserstein GAN conditioned with auxiliary data. They use ‘soul samples’, samples representing the ‘ideal’ sample from a class. Soul samples for seen classes are the average sample over all the seen samples for that class, while soul samples for unseen classes are constructed using the auxiliary data for the class. The GAN is trained to generate data close to a soul sample. The authors perform a step where unseen data is tested and any samples with high confidence are injected into training. This violates the requirement in zero-shot learning where unseen data should not be seen during training, even if it is unlabelled. GDAN [37] proposes a GAN with a regressor and cyclic-consistent loss. The results are marginally better than the conditional VAE method from Mishra et al. [34], indicating that GAN methods still fall behind VAE methods for zero-shot learning, given that at this point better VAE methods had been introduced.
Task Aligned Generative Meta-learning for Zero-shot Learning [38] is a proposed method to reduce bias towards particular classes by aligning the distribution of seen and unseen classes. The backbone of the technique is a meta conditional GAN [64]. SAGAN [39] involves training a GAN, followed by training a classifier to align the visual and semantic space. This method considers what to do with the trained GAN rather than improving the GAN training method. In the same year, ZeroNAS [65] is proposed which performs neural architecture search on GAN zero-shot learning methods to find the optimal architecture for a zero-shot learning problem. The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of the expert curation of these networks.
While there is much literature on how GANs may be used for zero-shot learning and how general improvements to GAN training can be used, they are still notoriously hard to train [63, 65]. This does pose the question of if GANs are too unstable and not reliable enough for zero-shot learning problems. In practical situations, it is hard to estimate model performance on unseen classes. Therefore, the method must be reliable.
Chou et al. [35] propose a generative method without VAEs or GANs. It uses a beta distribution to generate virtual classes for more robust zero-shot learning. Despite their efforts, GANs and VAEs still dominate the landscape of zero-shot learning.
In 2023, two publications emerged using diffusion in zero-shot learning [57, 58]. These both exist within the image classification domain and use the pre-trained Stable Diffusion model [47]. These provided insight into the power of diffusion models for zero-shot learning; however, the pre-trained nature of the model and the unknown training set make it hard to measure the ability of diffusion to perform zero-shot learning.
II-C Environmental Audio Zero-Shot Learning
Limited research exists for environmental audio zero-shot learning. One of the earliest works proposes a Twin neural network [66], tested on a small dataset containing ten classes. Xie and Viranen [18] used a more substantial dataset of fifty classes. In addition to a bilinear compatibility model [67] that created a baseline for the ESC-50 dataset, they proposed two partition strategies for five-fold cross-validation. These partitions are used in subsequent works in the field to allow researchers to compare methods reliably.
Further work in the field [68, 69] brought new state-of-the-art results and critical considerations when training zero-shot networks. Xie and Viranen [68] highlight the importance of the unseen and seen class distinction, where they identify that VGGish pre-trained on AudioSet is unsuitable for zero-shot learning on ESC-50 due to the presence of unseen classes in AudioSet.
AudioSet contains embeddings rather than audio clips. Embeddings are a critical aspect of zero-shot learning methods, and some research works explore improving these embeddings [19]. Furthermore, it is vital to maintain the distinction between ‘seen’ and ‘unseen’ classes. Therefore, training from raw data is essential. AudioSet provides the location of the YouTube videos the clips come from, but it is against the Terms of Service to download the raw data. AudioSet is, therefore, unsuitable for zero-shot learning research.
Some research focuses on improvements through embeddings. While not often the main focus of methods, they are the foundations of the zero-shot method. The remaining pipeline is limited to the information that the embeddings can capture. The methods in this paper use work by Sims et al. [19] to improve the performance through embeddings by adding synonyms.
III Method
This section will begin with a technical overview of zero-shot learning to complement the high-level background given in the previous section. Following this background information, the novel zero-shot learning diffusion method (ZeroDiffusion) is presented. A visual description of the method is shown in Figure 2.
III-A Zero-Shot Learning
The zero-shot learning problem is defined as the following. There exists a set $X$ of seen data and a set $Y$ of unseen data such that $X\cap Y=\varnothing$. The zero-shot learning method aims to learn a function
$f(z,C)=c$ (1)
where $z\in X\cup Y$ is a data sample; $C$ is the set of all classes; and $c\in C$ is the class that the sample $z$ comes from. After training a model to learn $f$ using all samples in $X$, the model is tested on all samples in $Y$.
The literature frequently divides zero-shot learning methods into three components. One component is the auxiliary data function $\sigma(c)$ that takes a class $c\in C$. The literature overwhelmingly uses word embedding models or human-engineered attributes for $\sigma$. $\sigma$ changes a class label into high-level information.
A second component is the feature embedding function $\tau(z)$, where $z\in X\cup Y$ is any feature data in the dataset. The feature data in the case of environment audio is audio clips. $\tau$ will convert the feature data into high-level information, similarly to $\sigma$. The leading method for $\tau$ in the literature involves training a traditional supervised classification model on the seen classes $Y$ and taking the second last layer as an information-rich embedding layer. In computer vision, a popular method is to use ResNet, pre-trained on ImageNet.
The third component ties the previous two together to retrieve a class prediction. This component learns a function
$\xi(\tau(z),\sigma(C))=c$ (2)
that takes the high-level information output of the other two components and outputs the corresponding class $c$ for a sample $z$. This component of zero-shot learning is explored heavily across the literature.
III-B Embeddings
An embedding is an encoded form of data that is easy for a machine learning model to process. They are typically information-rich low-dimensional vectors representing data that would be cumbersome and computationally intensive in their raw form.
In the context of zero-shot learning, embeddings are beneficial when they contain semantic information about the data. Learning the relationship between the feature and class semantic embeddings is crucial to performing zero-shot learning.
The $\sigma$ function is the Word2Vec [60] pre-trained model. This model takes class labels and outputs an embedding vector for that word. The space that Word2Vec has learned is information-rich, with semantically similar words existing closer together in the space than semantically distant words. Word2Vec gives an embedding in $\mathbb{R}^{300}$.
We use work by Sims et al. [19] to enrich the word embeddings with synonyms. The average word embedding of the class label and synonyms are used as auxiliary information. For ESC-50, the synonyms from the original paper were used. FSC22 has not been used in previous literature on synonyms, and so we human-engineer the synonyms, where overlapping classes use the same synonym list from ESC-50.
To learn $\tau$ and obtain audio embeddings, we train a modified YAMNet model, shown to perform better than other methods [19], on the seen class data. As shown in Figure 2, the second-last layer of the model is a fully connected embedding layer of size 128, designed to capture high-level semantic information for use as an audio embedding vector.
Thus both $\sigma$ and $\tau$ have been defined, where the output of $\sigma(c)$ is a class embedding of dimension $300$ for the class $c$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w\in X\cup Y$.
III-C ZeroDiffusion
Figure 2: ZeroDiffusion involves training three models. The first is the audio embedding model, used to retrieve audio embeddings. Seen classes are used to train this model. Next, the diffusion model is trained using audio embeddings from the first model and word embeddings. This process is done using seen classes. Finally, a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
Diffusion involves training a model to reconstruct data from noise. During training, the input to the model is a data sample with noise added, and the expected output is the original data sample. The amount of noise added to the data increases as training progresses, where the first few epochs have very little noise added to the data and the final epochs result in mostly or entirely noise. This process gradually trains a model to generate data from noise.
Conditioning a diffusion model allows the user to control the output. Image generation tasks condition the diffusion model on textual data descriptions. This conditioning allows the user to give a trained diffusion model a description of any image they want to generate [70]. If a user requests output of ‘a cat holding a basketball’, the model should be able to recreate a picture of a cat holding a basketball from noise.
A diffusion model conditioned on auxiliary data is trained with the input to the model being paired embeddings $\sigma(c)$ and $\tau(w)$ where $c$ is the corresponding class for the audio sample $w\in X$. The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(c)$.
The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with mean 0 and variance 0.1 to create a gradually noisy embedding vector
$\tilde{\tau}(w)(p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
where $p$ is the epoch percentage.
The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(c)$. The first embedding has a dimension of 128, and the second has a dimension of 300, resulting in an input layer of size 428. Due to the concise and information-rich nature of embeddings, the diffusion architecture is small and uses fully connected layers. The first hidden layer is a fully connected layer of size 128, to reduce the dimension to the expected output size. This is followed by batch normalisation for training stabilisation and then a Leaky ReLU activation to add a non-linearity component. Following Leaky ReLU, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space. The network output is an embedding of size 128 to match the size of the expected recreated audio embedding.
The loss is the mean squared error between the embedding the model is trying to recreate and the output of the model. During training, gradients are clipped to a maximum norm of 1 to stabilise training. An Adam optimiser is used.
Our new diffusion model is trained on seen class data. The aim is for the model to generalise to generating unseen data. Unlimited synthetic data for the unseen classes can be generated using the unseen class embeddings and random noise. After creating an unseen dataset in this manner, a bilinear compatibility network [67, 18] is trained with both the true seen data and synthetic unseen data. The bilinear network uses a cross-entropy loss function and an Adam optimiser.
IV Experiments
Experiments are performed on the ESC-50 and FSC22 datasets. The state-of-the-art in zero-shot learning environmental audio, the compatibility bilinear method [67, 18] is compared with existing methods from computer vision [14, 15], and our new method ZeroDiffusion.
ESC-50 is split into five partitions of ten classes. One partition is held as the test set. The remaining four partitions are used for four-fold cross-validation, where three of the partitions are used in training and the remaining one is used for validation. The validation set is used to tune the model and its hyperparameters before using the test set. This design is modified from the five-fold cross validation used in previous work [68, 19] to include a held-out test set. Using a four-fold strategy for hyperparameter tuning allows for more generalised hyperparameter tuning and a better understanding of the performance of the model. The final experiments on the held-out test set give a good indication of how the model performs on a completely unseen set of classes and data. ESC-50 is a suitable dataset for this strategy as it is composed of many classes, allowing for each partition to have ten classes. The dataset is correctly labelled and clean, making it ideal for testing.
FSC22 is split into three sets, a test set with 7 classes, a validation set of 7 classes, and a training set with the remaining 13 classes. The validation set is used for hyperparameter tuning, and then the method is evaluated using the test set. This dataset has almost half the number of classes compared to ESC-50, reducing the amount of flexibility in forming a strategy. Additionally, the number of classes is uneven and has few divisors. If it were partitioned into nine lots of three, the number of classes in each set would be extremely small for zero-shot learning while also creating a lot of folds to manage. If three lots of nine were used, the training set would be very small compared to the usual split in zero-shot learning, where there are more training classes than test classes. Therefore, the validation and test sets are chosen to be the same size for consistency, and train with a larger number, being almost the size of the test and validation sets combined. Therefore, $k$-fold validation is not performed for this dataset. The dataset does overlap with the classes in ESC-50, but there are new classes not present in ESC-50. Some classes have questionable labels, including the class ‘silence’, which appears to contain a variety of soundscapes including birds chirping and rain falling, which are classes elsewhere in the dataset. Because the dataset is not entirely well-labelled, this is considered a harder dataset compared to ESC-50.
All experiments are repeated ten times with the average and standard deviation of the accuracy recorded. The experiments were run on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
Index Class Index Class Index Class
0 Dog 17 Pouring water 34 Can opening
1 Rooster 18 Toilet flush 35 Washing machine
2 Pig 19 Thunderstorm 36 Vacuum cleaner
3 Cow 20 Crying baby 37 Clock alarm
4 Frog 21 Sneezing 38 Clock tick
5 Cat 22 Clapping 39 Glass breaking
6 Hen 23 Breathing 40 Helicopter
7 Insects (flying) 24 Coughing 41 Chainsaw
8 Sheep 25 Footsteps 42 Siren
9 Crow 26 Laughing 43 Car horn
10 Rain 27 Brushing teeth 44 Engine
11 Sea waves 28 Snoring 45 Train
12 Crackling fire 29 Drinking, sipping 46 Church bells
13 Crickets 30 Door knock 47 Airplane
14 Chirping birds 31 Mouse click 48 Fireworks
15 Water drops 32 Keyboard typing 49 Hand saw
16 Wind 33 Door, wood creaks
TABLE I: Class indices and names in ESC-50. This follows the convention from the dataset.
Index Class Index Class Index Class
0 Fire 9 Axe 18 Footsteps
1 Rain 10 Chainsaw 19 Clapping
2 Thunderstorm 11 Generator 20 Insect
3 Water Drops 12 Handsaw 21 Frog
4 Wind 13 Firework 22 Bird Chirping
5 Silence 14 Gunshot 23 Wing Flapping
6 Tree Falling 15 Wood Chop 24 Lion
7 Helicopter 16 Whistling 25 Wolf Howl
8 Vehicle Engine 17 Speaking 26 Squirrel
TABLE II: Class indices and names in FSC22. While the same order is followed from the dataset, here the classes are zero-indexed rather than one-indexed.
Partition Classes
ESC-50 fold 0 2, 3, 27, 29, 31, 35, 38, 40, 46, 48
fold 1 13, 19, 21, 22, 26, 32, 36, 39, 42, 49
fold 2 4, 10, 14, 17, 23, 24, 30, 33, 41, 45
fold 3 1, 6, 7, 18, 20, 25, 28, 34, 44, 47
fold 4 0, 5, 8, 9, 11, 12, 15, 16, 37, 43
FSC22 train 0, 1, 2, 3, 4, 10, 11, 14, 16, 19, 20, 24, 25
val 6, 8, 9, 12, 13, 18, 22
test 5, 7, 15, 17, 21, 23, 26
TABLE III: Classes in each partition. For ESC-50, random partitions are used from previous work for consistency [18]. FSC22 has not been benchmarked for zero-shot learning before, so a new partition strategy is created with a random number generator.
A modified YAMNet model [19] is trained using only the seen classes in FSC22. A learning rate of $10^{-3}$ is used and early stopping occurred at epoch 259 with an accuracy of $0.6833$. This model is used to retrieve the audio embeddings. The word embedding model is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/. Each method uses the same embedding vectors for each respective class and data sample.
IV-A Bilinear Compatibility
The bilinear compatibility network has been used to great success in environmental audio [18, 68, 19]. It represents the state-of-the-art method in zero-shot learning for environmental audio on ESC-50. Therefore, it is important to include it in this comparison.
The bilinear compatibility method has few hyperparameters and the parameters used by Xie and Virtanen [18] for ESC-50 were repeated in our work. FSC22 was tuned to train for 30 epochs.
IV-B CADA-VAE
Cross and distribution aligned variational autoencoder (CADA-VAE) [14] is a successful zero-shot learning method in computer vision, and for this reason it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
Hyperparameter tuning was conducted by starting with the default parameters from the computer vision experiments and manually tuning based on the validation sets. Both datasets used the same hyperparameters, aside from the number of classifier epochs. Overall, hyperparameter tuning was more useful in stabilising training and reducing the standard deviation than improving the accuracy. It was found that the large amount of hyperparameters made this method difficult and time consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
The VAE is trained with warmup strategies for the Kullback–Leibler divergence (KLD), the cross-reconstruction error, and the distance error. The KLD has a factor of one with a warmup over epochs 0 to 93, the cross-reconstruction has a factor of 2.37 with a warmup over epochs 21 to 75, and the distance error has a factor of 5.13 with a warmup over epochs 6 to 22. The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs.. The classifier has a learning rate of $10^{-3}$, a batch size of 32 and is trained for 23 epochs for ESC-50 and 20 epochs for FSC22. The loss of the VAE is mean absolute error and the loss of the classifier is negative log likelihood loss. Both use an Adam optimiser. The size of the latent space is 64.
IV-C LisGAN
Leveraging invariant side generative adversarial network (LisGAN) [15] is a more recent GAN method that reports comparatively high results amongst GANs in zero-shot learning. While there are some methodological concerns, these were modified to adhere to the zero-shot learning problem formulation and the core idea of the method remained. A conditional Wasserstein GAN is the main GAN method found in zero-shot learning. However, despite efforts to improve the method, it is still notoriously difficult to train [71, 65]. Smaller experiments were performed with a basic conditional GAN with no success, and often results were no better than random with a high variance across runs. LisGAN showed to be significantly more stable and competitive amongst other zero-shot learning methods.
Hyperparameter tuning of LisGAN involved simple manual tuning, with only a small amount of parameters to tune. These mostly were modified to suit the smaller ESC-50 and FSC22 datasets, as the datasets in computer vision are significantly larger. Both datasets used the same hyperparameters.
LisGAN is used without retraining on unseen classes. The GAN is trained with a batch size of 64, 128 units in the hidden layer of the discriminator and generator, 50 epochs and a learning rate of $10^{-4}$. The classifier is trained with a learning rate of $10^{-2}$ and 50 epochs. The generator, discriminator and classifier use Adam optimisers. The classifier uses a negative log likelihood loss.
ESC-50 FSC22
Val Fold 0 Val Fold 1 Val Fold 2 Val Fold 3 Test Val Test
Bilinear $0.2460\pm 0.0074$ $0.5003\pm 0.0074$ $0.3345\pm 0.0105$ $0.2868\pm 0.0076$ $0.2257\pm 0.0105$ $0.3526\pm 0.0069$ $0.3343\pm 0.0169$
LisGAN $0.2605\pm 0.0227$ $0.4039\pm 0.0220$ $0.2582\pm 0.0187$ $0.2150\pm 0.0195$ $0.2448\pm 0.0165$ $0.299\pm 0.0219$ $0.2142\pm 0.0416$
CADA-VAE $\textbf{0.2915}\pm 0.0180$ $0.3980\pm 0.0181$ $0.2500\pm 0.0251$ $0.3013\pm 0.0251$ $0.2095\pm 0.0337$ $0.3179\pm 0.0137$ $0.3290\pm 0.0339$
ZeroDiffusion $0.2846\pm 0.0111$ $\textbf{0.5279}\pm 0.0283$ $\textbf{0.3593}\pm 0.0189$ $\textbf{0.3342}\pm 0.0224$ $\textbf{0.3091}\pm 0.0278$ $\textbf{0.3975}\pm 0.0132$ $\textbf{0.3467}\pm 0.0200$
TABLE IV: Average accuracy and standard deviation over ten randomly seeded runs for the traditional bilinear compatibility method and the three generative methods for ESC-50 and FSC22. Our method, ZeroDiffusion, outperforms the other methods in accuracy across the partitions of both datasets. The bilinear compatibility method has less variance than the three generative-based methods.
IV-D ZeroDiffusion
ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in this paper. These experiments aim to determine the performance of this new method in the context of environmental audio. The existing success of diffusion in other areas of machine learning research motivate the creation and experimentation of this method.
The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 2000 epochs. Fifty samples are generated from the diffusion model per unseen class. The bilinear compatibility network is trained with a learning rate of $10^{-4}$, batch size of 52, and 20 epochs. These hyperparameters were chosen using manual tuning.
V Results
The results from experiments with the four methods - bilinear compatibility, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50 and FSC22 are presented in Table IV.
ZeroDiffusion increases the accuracy on the ESC-50 and FSC22 test partition by $36.95\%$ and $3.71\%$, respectively, from the baseline bilinear method. With the exception of the first validation fold in ESC-50, ZeroDiffusion has a higher accuracy than the other three zero-shot learning methods. This shows that ZeroDiffusion is a new state-of-the-art method for both ESC-50 and FSC22.
Notably, ZeroDiffusion incorporates the bilinear model in its methodology. The difference between the methods is that ZeroDiffusion trains a diffusion model to generate synthetic unseen data and adds these samples to the training of the bilinear method. Because the diffusion model is the only addition, it is reasonable to conclude that the diffusion model is responsible for the improved accuracy.
CADA-VAE has positive performance but is inconsistent across partition strategies. LisGAN has similar inconsistency issues but overall appears to have worse performance across partitions and datasets. This inconsistency continues when considering the standard deviation for each experiment. The generative methods have high variance compared to the baseline method. In the case of ZeroDiffusion, the higher average accuracy and variance are tradeoffs to consider.
The accuracy improvement of ZeroDiffusion supports the use of diffusion models in the zero-shot learning field. Adjacent works with Stable Diffusion [57, 58] are supported by ZeroDiffusion’s performance, indicating that diffusion methods more broadly have potential in a range of domains for zero-shot learning.
Confusion matrices give more insight into model performance through the analysis of class predictions. The confusion matrices for one run of ESC-50 fold 0 and 1 are shown in Tables V and VI, respectively.
In fold 0 of the ESC-50 dataset, the helicopter and mouse click classes have the highest accuracy at 80% and 62.5%, respectively. The helicopter class is predicted at least 70% of the time for four different classes, resulting in a hub around the helicopter class. A hub is a class where predictions across classes are clustered, i.e. the model is much more likely to predict the hubs than other classes. In zero-shot learning, the hubness problem has been identified as a major issue preventing classes from being predicted, resulting in a low accuracy. In fold 0, smaller hubs exist around drinking, brushing teeth and mouse clicking. Other classes are not predicted often. The difference between the hub classes and other classes does not appear to have a clear intuitive explanation. This information could lead to avenues for improvement.
Fold 1 of ESC-50 has a significantly higher accuracy than fold 0. The highest class accuracy is siren at 80%, however all but the clapping class have at least 25% accuracy. Clapping is predicted a total of three times across all clips in this fold, and all three are misclassified hand saw clips. Other classes are predicted more consistently, indicating that the hubness problem isn’t as prevalent in the training of this partition of the dataset.
The confusion matrix in fold 1 is sparse, where classes are only misclassified as a small number of other classes. This phenomenon is also found in traditional supervised learning, where the target class is misclassified as classes with similar features. In zero-shot learning, it is the machine learning-based semantic embeddings that are being compared, and so intuition from traditional supervised learning may not apply.
Some misclassifications appear intuitive, such as the case where the sound of a washing machine is incorrectly labeled as a helicopter. Both classes are loud and mechanical, motivating a potential misunderstanding from the model. Fireworks are misclassified as mouse clicking in fold 0, and the reason may not initially be intuitive, but both sounds are comprised of short, sharp noises. There is a difference in loudness, but the audio data is normalised before input to the network and so this difference may be reduced. While there are some intuitive misclassifications, there are also non-intuitive misclassifications. Pig is incorrectly labeled as brushing teeth, and door knocking is mislabeled as rain.
To gain a deeper understanding, dimensionality reduction with t-distributed stochastic neighbour embedding (t-SNE) [72] on the audio embeddings produces Figure 3 for fold 0 and Figure 4 for fold 1. In fold 0, pig (class 2, dark purple) is spread across a wide area, including overlapping with brushing teeth (class 27, dark blue). This relationship gives insight into the non-intuitive misclassification of pig.
Fold 1 has a significantly higher accuracy than fold 0, as shown in Table IV. The t-SNE graph for fold 1 shows more distinct clusters than fold 0, indicating that this clustering property of the audio embeddings is important in correctly classifying unseen classes. This is reinforced when considering the clapping class, which both is the poorest performing class in fold 1 and has the greatest spread across the t-SNE graph.
The t-SNE graphs reveal insights into the hubness problem, and asymmetry in the incorrect predictions further highlights the problem. Pig is mislabeled as brushing teeth more than 50% of the time, while brushing teeth is never mislabeled as pig. This shows that certain classes are more likely to be predicted than others. The classes in fold 0 that are frequently predicted as the helicopter class are closer together in the t-SNE graph, showing their similarity in the embedding space. This similarity is a potential cause of the hubness problem.
Class index 38 29 35 3 40 2 27 46 31 48
38 1 9 0 3 0 3 9 0 14 1
29 1 18 0 0 1 0 18 0 1 1
35 1 0 6 0 30 1 1 1 0 0
3 0 0 5 0 29 5 1 0 0 0
40 0 0 6 0 32 0 0 1 0 1
2 0 6 3 0 3 5 21 0 0 2
27 2 34 0 1 0 0 1 0 2 0
46 0 0 2 0 28 2 0 8 0 0
31 1 5 0 0 2 0 6 0 25 1
48 4 6 0 0 2 0 8 1 15 4
TABLE V: Confusion Matrix for fold 0 of the ESC-50 dataset. Rows represent the true class, and columns represent the predictions. Many predictions are clustered around class 40, with smaller hubs on classes 29, 27 and 31.
Figure 3: ESC-50 fold 0 audio embeddings with t-SNE dimensionality reduction. Class 40 (helicopter), 46 (church bells), 35 (washing machine) and 3 (cow) are all concentrated to the right side of the scatterplot. The confusion matrix in Table V shows that all of these classes are classified as helicopter a majority of the time.
Class index 39 36 42 13 32 22 19 49 26 21
39 10 0 0 0 9 0 2 3 0 16
36 0 24 3 4 0 0 6 1 0 2
42 0 0 32 3 0 0 0 2 3 0
13 0 4 7 14 1 0 5 0 1 8
32 2 0 0 0 23 0 2 13 0 0
22 14 1 0 1 3 0 14 6 1 0
19 0 5 10 0 0 0 22 0 0 3
49 1 3 0 1 0 3 0 11 8 13
26 4 0 2 4 0 0 0 4 15 11
21 2 0 0 1 1 0 0 0 8 28
TABLE VI: Confusion matrix for fold 1 of the ESC-50 dataset. Rows represent the true class, and columns represent the predictions. There are small hubs on classes 42, 19, 26 and 21 and all classes are at least partially correctly identified, except for class 22 which is never correctly classified.
Figure 4: ESC-50 fold 1 audio embeddings with t-SNE dimensionality reduction. Fold 1 exhibits more defined class clusters than fold 0. Class 22 (clapping, dark blue) has a greater spread, and is the poorest classified class as shown in Table VI.
t-SNE graphs give a deeper insight into issues surrounding zero-shot learning, and can indicate avenues for improvement. The relationship between the clustering properties of the audio embeddings and model accuracy show that embeddings have an effect on the performance of a zero-shot learning model. More defined clustering in the audio embeddings produces higher accuracy, and hubs in the confusion matrix are reflected in the t-SNE diagram. The confusion matrices highlight the abilities of the model beyond an accuracy value. The sparsity of the matrices demonstrate that the model recognises the similarity of clips from the same class, and when it correctly maps these to the correct class the performance is very high.
VI Discussion
In a true zero-shot learning real-world application, it is hard to know how well a model will perform on the target class due to a lack of available samples. Therefore, consistency across training runs is important. While ZeroDiffusion has a superior accuracy-based performance, the bilinear method has superior consistency. This is an important consideration when choosing an appropriate method for zero-shot learning in practical applications.
The hubness problem is an issue when considering practical usage because of the poor accuracy on specific individual classes that are overshadowed by the hubs. Many of these specific classes are predicted at a rate worse than a random model. Hubness appears to correlate to overall model performance, as seen when comparing the confusion matrices and overall accuracy of two partitions of the ESC-50 dataset.
One key avenue in reducing the hubness problem is first understanding why the different partitions used in this work had varying levels of hubness. Partitions can be crafted for this purpose to identify the effect of different criteria on hubness. One question is whether the unseen classes, seen classes, or both, influence the performance. If particular characteristics in the seen class set influence the hubness problem, then the seen class set can be supplemented with classes exhibiting these characteristics. A more likely scenario is that both influence the performance, and the seen classes used in training may need to be curated for the characteristics of the specific unseen set in the given problem.
Zero-shot learning lacks theoretical foundations, with methods developing through empirical results. There is little understanding of upper bounds to accuracy for a given unseen and seen class sets. A given seen class set will not necessarily have enough information to extrapolate to particular unseen classes. A trivial example is if the seen class set is the empty set. A deeper understanding of these requirements may be an avenue to reducing hubness, and therefore improving overall classification performance.
VII Conclusion
A comparison of three generative methods and the baseline method in environmental audio zero-shot learning has been presented. Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly or worse than the baseline, and present higher variance.
The third generative method, ZeroDiffusion, is a novel diffusion-based method for zero-shot learning. It shows a higher accuracy performance than the other three methods, thus producing a new state-of-the-art result in the domain of environmental audio on ESC-50 and FSC22. Diffusion shows potential in zero-shot learning to improve the performance across domains. Future work could include benchmarks for diffusion-based zero-shot learning models on datasets in areas other than environmental audio.
An analysis of confusion matrices and t-SNE graphs of the audio embeddings reveals the impact of audio embedding similarities on the hubness problem and model performance. This analysis is valuable since these similarities may not be intuitive from the raw audio data.
Acknowledgements
This research is supported by an Australian Government Research Training Program Scholarship to the first author.