### v2

# Embedding-Space Diffusion for Zero-Shot Environmental Sound Classification

Abstract
Zero-shot learning enables models to generalise to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes. While much research has focused on zero-shot learning in computer vision, the application of these methods to environmental audio remains underexplored, with poor performance in existing studies. Generative methods, which have demonstrated success in computer vision, are notably absent from zero-shot environmental sound classification studies.
To address this gap, this work investigates generative methods for zero-shot learning in environmental audio. Two successful generative models from computer vision are adapted: a cross-aligned and distribution-aligned variational autoencoder (CADA-VAE) and a leveraging invariant side generative adversarial network (LisGAN). Additionally, we introduced a novel diffusion model conditioned on class auxiliary data. Synthetic embeddings generated by the diffusion model are combined with seen class embeddings to train a classifier.
Experiments are conducted on five environmental audio datasets, ESC-50, ARCA23K-FSD, FSC22, UrbanSound8k and TAU Urban Acoustics 2019, and one music classification dataset, GTZAN. Results show that the diffusion model outperforms all baseline methods on average across six audio datasets.
This work establishes the diffusion model as a promising approach for zero-shot learning and introduces the first benchmark of generative methods for zero-shot environmental sound classification, providing a foundation for future research.
Index Terms:
Environmental Audio, Zero-Shot Learning, Generative Networks
I Introduction
Machine learning has revolutionised how society solves complex problems [2, 3, 4]. Various publications consider the future impact of machine learning, such as smart cities [5], humanoid robotics [6] and medicine [7, 8, 9]. To solve these future problems, machines and AI need to be adaptive [10, 11]. One way machines can be adaptive is to classify new items without labelled training data [12, 13]. One method to achieve this is zero-shot learning.
Zero-shot learning is a subset of supervised machine learning that aims to create a model to classify classes it has not seen during training. That is, the set of classes used in training and testing are disjoint. Zero-shot learning is valuable in real-world applications, such as remote sensing [14], image classification [15, 16], image segmentation [17, 18], environmental audio [19, 20] and security [21], where there may not exist data for every class.
Outside of zero-shot learning, recent practical use cases for environmental audio classification include smart cities [22], detection of traffic noise pollution [23], and hearing aid analysis [24]. These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
An intuitive way to understand zero-shot learning involves a thought experiment on how humans would solve the problem. Suppose someone has never seen a snake, cat or bird but knows key attributes these animals have. When shown a picture of a cat, the person reasons that it is a cat as it has four legs and whiskers, and no scales or wings. The person performs zero-shot learning using attributes of the target class and learning what those attributes mean through knowledge of other animals.
The presented work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision. It takes inspiration from the successful usage of generative methods in computer vision [15, 25, 26, 27], and from a newer generative method, diffusion, which has seen success in image generation tasks [28, 29]. There is currently no literature known to the authors on environmental audio zero-shot learning using generative methods.
Six datasets - ESC-50, FSC22, ARCA23K-FSD, TAU Urban Acoustic Scenes 2019, UrbanSound8k and GTZAN - are used to compare four methods - the baseline non-linear compatibility network [30] based on Attribute Label Embedding (ALE) [31], LisGAN [16], CADA-VAE [15] and our new novel method, ZeroDiffusion. This represents a broad study of generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion for zero-shot learning, with the baseline as a comparison. ZeroDiffusion achieves significant accuracy improvement in most cases.
The contributions of the presented paper are:
  • The first diffusion method for zero-shot learning that does not use a pre-trained model, but rather works in the embedding space.
  • A new state-of-the-art method for zero-shot environmental sound classification that performs equally to or outperforms a range of existing methods in zero-shot learning across six audio datasets.
  • The first zero-shot learning benchmark on the FSC22, ARCA23K-FSD and TAU Urban Acoustic Scenes 2019 datasets.
  • The first comparison of GANs, VAEs and diffusion-based methods for zero-shot environmental sound classification with six audio datasets: ESC-50, FSC22, ARCA23K-FSD, UrbanSound8k, TAU Urban Acoustic Scenes 2019 and GTZAN.
II Related Work
This section first covers terminology in zero-shot learning to better understand the literature. Then, it examines existing work on generative methods in zero-shot learning, much of which is in the domain of computer vision. Next, existing works in zero-shot environmental audio classification are considered and audio datasets are covered. Finally, with this context, the novelty and contributions of the presented work are explained.
II-A Terminology
To contextualise the literature, this subsection introduces foundational terminology in zero-shot learning. It first describes class and feature embeddings, followed by the role of auxiliary data. A high-level definition of zero-shot learning is then provided to frame the broader research landscape.
II-A1 Embeddings
Embeddings convert raw data, such as images or audio, into a high-level form that machine learning algorithms can use. The form embeddings take is a high-dimensional vector of real numbers, formally $x\in\mathbb{R}^{\alpha}$ is an example of an embedding with dimension $\alpha$. The methods in this work use deep neural networks to obtain these vector representations, using both supervised and unsupervised methods. An example of what an embedding space might look like is shown in Figure 1. Lions and cats are semantically similar and, therefore, are closer to each other in the vector space than a dog or car. In this way, the structure of the space reveals semantic relationships.
Figure 1: An embedding space. Semantically similar data is closer, such as ‘lion’ and ‘cat’.
Supervised methods to create embeddings typically involve classical supervised learning on a dataset, and an internal layer of the artificial neural network becomes the embedded data vector. In many cases, the second last layer is used because it is expected to contain high-level meaningful information compared to previous layers.
Zero-shot learning requires a clear distinction between the seen and unseen classes, and so the dataset used to learn the embedding must only contain classes that are not part of the unseen class set. Typically, the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset is suitable as long as there is no overlap with the unseen class set.
Unsupervised methods involve learning a mapping to a space that represents the semantic distance of data. A common example of this type of embedding method is word embeddings, such as GloVe [32] and Word2Vec [33]. These methods learn the semantic similarity of words using their proximity in texts. To a human, embeddings may seem to contain little meaning, but these vectors are information-rich. For example, the word embedding vectors for ‘man’, ‘woman’ and ‘king’ can be combined in an equation ‘king’ - ‘man’ + ‘woman’ to approximate the vector for ‘queen’ [33]. GloVe uses matrix factorisation and builds a large co-occurrence matrix of word pairs to globally understand the relationship between words, while Word2Vec uses context-based predictive neural networks. While their learning processes are different, both methods have the same goal and provide semantically rich embeddings.
II-A2 Auxiliary Data
Zero-shot learning involves learning to classify known classes without samples for those classes. If the classes are unknown, the problem is open-set learning. These known classes contain information even without samples. This information is ‘auxiliary data’. The most common form of auxiliary data is the word embedding of the class label, retrieved using Word2Vec or GloVe. Another popular form is human-engineered attributes, where a vector is manually created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’. The Animals with Attributes [13] and Animals with Attributes 2 [34] datasets are examples of datasets that contain human-engineered attribute vectors and were created for zero-shot learning research.
II-A3 Zero-Shot Learning
Zero-shot learning involves training a model on a set of seen classes $X$ and evaluating its performance on unseen classes $Y$, where $X\cap Y=\varnothing$.
Feature data refers to the input the model is classifying, such as images in computer vision or audio clips in environmental sound classification. Feature data from unseen classes must not be used before testing, including pre-training.
Both feature data and class labels are represented as embeddings that capture semantic information. The goal of zero-shot learning is to train a model that can leverage these semantic representations to generalise beyond the classes it has been explicitly trained on.
II-B Generative Methods in Zero-Shot Learning
While there is no literature known to the authors in environmental audio that considers generative networks, there are many examples in the domain of computer vision. These largely involve networks combining GANs and autoencoders.
II-B1 Autoencoders
In 2017, autoencoders entered the computer vision zero-shot learning space through two publications, Tsai et al. [35] and Mukherjee et al. [36]. They used two traditional autoencoders for each modality (class textual descriptions and images) and trained a shared latent space between these two modalities. Conditional variational autoencoders were introduced in zero-shot learning the following year [26]. A new benchmark was found when CADA-VAE [15] was proposed, which used variational autoencoders and strengthened the shared latent space through cross-alignment and distribution-alignment of the two modalities.
II-B2 Generative Adversarial Networks
One of the first GAN methods in zero-shot learning is f-CLSWGAN, a Wasserstein GAN method [37]. A Wasserstein GAN is a vanilla GAN with the loss function changed to measure the Wasserstein distance between the real and generated data distributions [38]. It results in more stable training and mitigates mode collapse. In the same year, Zhu et al. introduced another GAN method using a visual pivot regulariser to guide training [39]. The regulariser encourages the generation of features for a class to be close to a fixed pivot point. It used Wikipedia text descriptions converted to word embeddings using the term frequency-inverse document frequency (TF-IDF) as class data. The authors compare their method with other zero-shot learning methods, including a basic GAN model. The basic GAN model has much worse accuracy than any established zero-shot learning method on the vision classification task across benchmark datasets.
LisGAN [16] improved on previous GAN methods by using a Wasserstein GAN conditioned on auxiliary data. They use ‘soul samples’, samples representing the ‘ideal’ sample from a class, to guide generation. The method includes a self-training step where high-confidence samples from the unseen classes are injected into the training process. This makes the approach transductive, since it assumes access to unlabelled unseen data, in contrast to the inductive setting commonly assumed in zero-shot learning where no unseen data is available. GDAN [27] proposes a GAN with a regressor and cyclic-consistent loss. The results are marginally better than the conditional VAE method from Mishra et al. [26]. Given the existence of superior VAE methods, this indicates that GAN methods still fall behind VAE methods for zero-shot learning.
Task Aligned Generative Meta-learning for Zero-shot Learning [40] is a proposed method to reduce bias towards particular classes by aligning the distribution of seen and unseen classes. The backbone of the technique is a meta conditional GAN [41], which combines a meta-learning network and a conditional GAN. SAGAN [42] involves training a GAN, followed by training a classifier to align the visual and semantic space. This method focuses on improving the classifier rather than the GAN. In the same year, ZeroNAS [43] is proposed which performs neural architecture search on GAN-based zero-shot learning methods to find the optimal architecture for a zero-shot learning problem. The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of expert curation of these networks.
While there is much literature on using GANs in zero-shot learning, they are still notoriously hard to train [39, 43]. GANs may be too unstable and unreliable for zero-shot learning problems. In practical situations, it is hard to estimate model performance on unseen classes. Therefore, the method must be reliable.
II-B3 Diffusion
In 2023, two publications emerged using diffusion in zero-shot learning [44, 45]. These both exist within the image classification domain and use the pre-trained Stable Diffusion model [29]. These provided insight into the power of diffusion models for zero-shot learning; however, the pre-trained nature of the model and the unknown training set make it hard to measure the ability of diffusion to perform zero-shot learning. Additionally, since Stable Diffusion is foundational to the method, usage is limited to problems in computer vision and the use of textual class auxiliary data.
II-C Environmental Audio Zero-Shot Learning
Limited research exists for environmental audio zero-shot learning. An early study proposes a Twin neural network [46], tested on a small dataset containing ten classes. A Twin neural network is an architecture with two identical networks that process two inputs at the same time to learn a similarity function. Xie and Virtanen [19] used a more substantial dataset of fifty classes. They introduce a bilinear compatibility network, which builds on from Attribute Label Embedding (ALE) [31]. A compatibility function is learnt between the audio data and class embeddings by computing:
$f(x,y)=x^{T}Wy$ (1)
Where $x$ is the audio embedding, $y$ is the class embedding, and $W$ is a trainable weight matrix. This maps the inputs to a shared latent space for comparison directly, without the use of generative methods.
This method sets an initial baseline for the ESC-50 dataset, and the authors introduce two partition strategies for five-fold cross-validation. These partitions are used in subsequent works in the field to allow researchers to compare methods reliably.
Further work in the field [47, 48] brought new state-of-the-art results and critical considerations when training zero-shot networks. Xie et al. [47] highlight the importance of the unseen and seen class distinction when using pre-trained models. A VGGish model trained on AudioSet (a large environmental audio dataset) may produce rich embeddings, but a majority of classes in ESC-50 occur in AudioSet. The classes in ESC-50 which are meant to be unseen would have instead been seen by the VGGish network during pretraining on AudioSet. This breaks the zero-shot learning formulation. It is important to understand that it is the feature data that should be unseen - the classes are known, and therefore a pre-trained word embedding model does not invalidate the method as zero-shot learning.
Wu et al. use contrastive audio-text representation learning with CLAP to perform zero-shot learning on multiple classes[49]. However, like other CLAP and CLIP-based methods, these models require training data from large datasets in the target modality and this is not always feasible in a given domain. Additionally, even though repeated samples across datasets are removed, whole classes aren’t removed. This is evident in the ESC-50 and FSD50k case, where the text reports 630 overlapping files, even though within one overlapping class, ‘dog’, there are over 900 samples in FSD50K that would need to be removed. Indeed, this would result in extremely few classes remaining since audio sources for datasets are very limited and many datasets are from YouTube or Freesound, with many of the same classes appearing across audio datasets. However, this inflates the reported zero-shot classification performance.
Some research focuses on improvements through embeddings. While not often the main focus of methods, they provide the foundations for unseen classification. Any embedding-based method is limited to the information that the embeddings can capture. The methods in this paper use work by Sims et al. [20] to improve the performance through embeddings by adding synonyms, except ARCA23K-FSD (due to size) and GTZAN (lack of synonyms for genres).
II-D Contributions and Novelty
This section will detail the contributions of the presented paper, in the context of the literature. Literature for zero-shot environmental sound classification is lacking, and no existing literature in the domain explores generative methods. The experiments in this work not only benchmark environmental audio datasets on the existing generative methods CADA-VAE and LisGAN but also introduce a novel diffusion-based method that is not tied to a particular modality. Code for the novel ZeroDiffusion method is provided at https://github.com/ysims/ZeroDiffusion.
Two publications exist in zero-shot learning using diffusion ([44, 45]); however, they both use Stable Diffusion directly to classify unseen data, with no further training of any model. Their methods are specific to computer vision and textual class labels through their use of Stable Diffusion. Meanwhile, the novel diffusion method presented in this work is modality-agnostic, because it uses embeddings as input to the diffusion model. Any class embedding can be used to condition the diffusion model, and any feature embedding can be used for training, regardless of what modality it comes from. The method involves training a diffusion model, which is used to generate unseen embeddings to train a compatibility-based classifier. This generative-style zero-shot learning method is significantly different to the previous work on Stable Diffusion and provides greater flexibility.
In addition to designing the first modality-agnostic diffusion method for zero-shot learning, this work additionally performs the first zero-shot benchmarks of FSC22, ARCA23K-FSD and TAU Urban Acoustic Scenes 2019.
III Methods
This section will begin with a technical overview of zero-shot learning to complement the high-level background given in the previous section. Following this background information, the novel zero-shot learning diffusion method (ZeroDiffusion) is presented. A visual description of the method is shown in Figure 2.
III-A Zero-Shot Learning
The zero-shot learning method aims to learn a similarity function $f(w,z)$, where $w$ is a feature data sample and $z\in X\cup Y$ is a class. $X$ is the set of unseen classes and $Y$ is the set of seen classes. $X\cup Y$ is the set of all classes and $X\cap Y=\varnothing$. After training a model to learn $f(w,z)$ using all samples belonging to classes in $X$, the model is tested on all samples belonging to classes in $Y$.
Many zero-shot learning methods can be considered to contain three distinct components. One component is the auxiliary data function $\sigma(z)$ that takes a class $z\in X\cup Y$. The literature overwhelmingly uses word embedding models or human-engineered attributes for $\sigma$. $\sigma$ changes a class label into high-level information.
A second component is the feature embedding function $\tau(w)$, where $w$ is any feature data in the dataset. The feature data in the case of environment audio is audio clips. $\tau$ will convert the feature data into high-level information, similarly to $\sigma$. The leading method for $\tau$ in the literature involves training a traditional supervised classification model on the seen classes $Y$ and taking the second last layer as an information-rich embedding layer.
The third component ties the previous two together to retrieve a class prediction. This component learns a similarity function $\xi(\tau(w),\sigma(z))$ that takes the high-level output of the two embedding components and outputs the similarity of $w$ and $z$. By comparing $w$ against all classes ($X$ during training and $Y$ during testing), either top-1 or thresholding can be used to retrieve the prediction.
III-B Embeddings
An embedding is an encoded form of data that is simple for a machine learning model to process. They are typically information-rich low-dimensional vectors representing data that would be cumbersome and computationally intensive in their raw form.
In the context of zero-shot learning, embeddings are beneficial when they contain semantic information about the data. Learning the relationship between the feature and class semantic embeddings is crucial to performing zero-shot learning.
The $\sigma(z)$ function in this work is the Word2Vec [33] pre-trained model. This model takes class labels and outputs an embedding vector for that word. The space that Word2Vec has learned is information-rich, with semantically similar words existing closer together in the space than semantically distant words. Word2Vec gives an embedding in $\mathbb{R}^{300}$.
We use work by Sims et al. [20] to enrich the word embeddings with synonyms. The average word embedding of the class label and synonyms are used as auxiliary information. For ESC-50, the synonyms from the original paper were used. Synonyms are human-engineered for the other datasets, where overlapping classes across datasets use the same synonym lists. ACRA23K-FSD and GTZAN are used without added synonyms, due to size and difficulty in getting synonyms for music genres. Synonyms used are detailed in the supplementary materials.
To learn $\tau(w)$ and obtain audio embeddings, we train a modified YAMNet model [20] on the seen class data, due to the higher accuracy compared to VGGish when used for zero-shot learning. The input to the modified YAMNet model is a log Mel spectrogram with 64 bins. As shown in Figure 2, the second-last layer of the model is a fully connected embedding layer of size 128, designed to capture high-level semantic information for use as an audio embedding vector.
Both $\sigma(z)$ and $\tau(w)$ have been defined, where the output of $\sigma(z)$ is a class embedding of dimension $300$ for the class $z\in X\cup Y$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w$.
III-C ZeroDiffusion
Figure 2: ZeroDiffusion involves training three models. The first (1) is the audio embedding model, used to retrieve audio embeddings. The model is trained with seen classes. Next, (2) the diffusion model is trained using audio embeddings and word embeddings from seen classes. Finally, (3) a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
We introduce our new method ZeroDiffusion, which applies a novel embedding-space diffusion model to generate synthetic unseen data to train a compatibility classifier. While other diffusion methods use large diffusion models to generate feature embeddings, ZeroDiffusion uses feature embeddings as input to learn to generate unseen data. The diffusion model itself is very lightweight because it works within the embedding space.
Diffusion involves training a model to reconstruct data from noise. During training, the model input is a data sample with added noise, and the expected output is the original data sample. The amount of noise added to the data increases as training progresses, where the first few epochs have very little noise added to the data and the final epochs result in mostly or entirely noise. This process gradually trains a model to generate data from noise.
Conditioning a diffusion model allows the user to control the output. Image generation tasks condition the diffusion model on textual data descriptions. This conditioning allows the user to give a trained diffusion model a description of any image they want to generate [50]. If a user requests output of ‘a cat holding a basketball’, the model should be able to recreate a picture of a cat holding a basketball from noise.
A diffusion model conditioned on auxiliary data is trained where the input to the model is the paired embeddings $\sigma(z)$ and $\tau(w)$ where $z$ is the corresponding class for the audio sample $w$. The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(z)$.
The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with a mean of 0 and a variance of 0.1 to create a gradually noisy embedding vector:
$\tilde{\tau}(w,p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (2)
where $p$ is the epoch percentage. Similarly the class embedding $\sigma(z)$ is jittered for diversity using
$\tilde{\sigma}(z)=\sigma(z)+\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(z)$. The first embedding has a dimension of 128, and the second has a dimension of 300, resulting in an input layer of size 428. Due to the concise and information-rich nature of embeddings, the diffusion architecture is small and uses fully connected layers. The first hidden layer is a fully connected layer of size 128, to reduce the dimension to the expected output size. This is followed by a Leaky ReLU activation for a non-linearity component and then a dropout of $0.3$ to encourage diversity. Following the dropout layer, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space. The network output is an embedding of size 128 to match the size of the expected recreated audio embedding.
The loss term involves five components: a mean squared error for the reconstruction loss, a maximum mean discrepancy for distribution loss, variance loss to encourage diversity, an average mean squared error of generated and real features in a batch for global centroid alignment, and a cosine similarity loss to encourage directional alignment of embeddings. During training, gradients are clipped to a maximum norm of 1 to stabilise training. An Adam optimiser is used.
Our new diffusion model is trained on seen class data. The aim is for the model to generalise to generating unseen data. Unlimited synthetic data for the unseen classes can be generated using the unseen class embeddings and random noise. After creating an unseen dataset in this manner, a non-linear compatibility network [31, 48] is trained with both the true seen data and synthetic unseen data. The compatibility network uses a cross-entropy loss function and an Adam optimiser.
III-D Non-linear ALE
Xie et al. introduce the non-linear version [48] of the bilinear compatibility network [19], which is based on ALE [31]. This method is used as a baseline due to its good performance amongst the zero-shot environmental sound classification literature and its simplicity. The lack of generators (as opposed to GAN and VAE methods) and its single classification nature means that it has few hyperparameters to tune and training across runs is stable, as shown in the experiments presented in section IV.
The method learns a function:
$f(w,z)=(B\cdot(\tanh(A\cdot w)))^{T}\cdot z$ (4)
where $w$ is the audio embedding, $z$ is the class embedding, $A$ and $B$ are both trainable weight matrices, and $\tanh$ is the hyperbolic tangent activation. $f(w,z)$ represents the similarity between $w$ and $z$, and is calculated for every $z$. This creates an encoded vector where the $i$th position in the vector represents the confidence or similarity for the $i$th class. The model is independent of the number of classes, unlike typical classification models where the last layer is equal to the number of classes.
III-E LisGAN
LisGAN [16] is a GAN method that uses ‘soul samples’ to regulate data generation. It trains a conditional Wasserstein GAN [38] where the generator uses class embeddings and noise to generate feature data and the discriminator not only determines whether a sample is generated or real but also has a supervised classification component to guide the conditional training. Each class has a soul sample - a representative sample for the class - and the generator is trained to generate samples for a given class close to that class’s soul sample. For seen classes, the soul sample is the average of that class’s data. For unseen classes, the class embedding defines the soul sample.
The discriminator’s classification output is used to screen synthetic samples. Synthetic samples with high confidence are used to train a final classifier. Additionally, unseen and unlabelled data that the discriminator’s classifier predicts with high confidence are used in training the final classifier. This use of unlabelled unseen data is transductive zero-shot learning; however, in this work, only the inductive case is considered. Therefore, we remove the training on unlabelled seen data. The core idea of the LisGAN method remains, where soul samples are used to guide GAN training, and the discriminator’s classifier is used to filter data to train a final classifier. Despite removing a component of LisGAN, it outperformed other GAN methods (vanilla conditional Wasserstein GAN, f-CLSWGAN) in pilot studies.
III-F CADA-VAE
CADA-VAE [15] is a popular method in computer vision-based zero-shot learning that competes with new methods. In this work, CADA-VAE is applied to the domain of environmental audio to evaluate its performance as a variational autoencoder-based method for zero-shot environmental sound classification. The core of the method is the two variational autoencoders (VAEs) and the loss function. One VAE is designed to encode and decode class embeddings, while the other is designed to encode and decode feature embeddings (typically image embeddings, but in this work, audio embeddings are used). Not only are the VAEs trained in a typical manner with reconstruction loss and Kullback–Leibler divergence, but they are also trained with cross-alignment and distribution alignment. Cross-alignment measures the ability of the decoders to decode data from the other modality into their modality - that is, take an encoded class embedding and decode it into a feature embedding, and vice versa. Distribution alignment measures the difference between the distribution of the class embedding autoencoder and the feature embedding autoencoder. Both of these loss terms aim to create a shared latent space between both modalities. The latent space is then used to train a classifier, which performs zero-shot learning at test time.
IV Experiments
Experiments are performed on the ESC-50, ARCA23K-FSD, FSC22, UrbanSound8K, TAU Urban Acoustic Scenes 2019 and GTZAN datasets. The state-of-the-art in zero-shot learning environmental audio, the non-linear ALE method [31, 48] is compared with existing methods from computer vision [15, 16], and our new method ZeroDiffusion.
All experiments are repeated ten times with the average and standard deviation of the accuracy recorded. The experiments were conducted on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
IV-A Datasets
Environmental sound classification severely lacks large datasets and diversity. ESC-50, FSC22, ARCA23K-FSD and UrbanSound8k are sourced from Freesound, a repository of user-contributed tagged audio clips released under the Creative Commons license. Some clips in these datasets may overlap, and classes are typically similar. TAU 2019 and GTZAN are collected from other sources, and investigate the problems of acoustic scene recognition and music genre classification, respectively. These datasets add diversity to the experiments both in audio clip source and task.
IV-A1 ESC-50
ESC-50 [51] is a clean environmental sound classification dataset of 50 classes which can be arranged into five balanced categories with ten classes each. There are 40 audio clips per class, each five seconds long. The dataset is from Freesound and has been carefully curated to ensure a clean dataset. It is a popular dataset due to its reasonable size and correct labels. The large number of classes in the dataset motivates using k-fold cross-validation, as done in previous work [19, 20].
IV-A2 FSC22
FSC22 [52] is a sound classification dataset focusing on forest sounds. It has 27 classes, each with 75 5-second-long audio clips. The clips are sampled from Freesound. The dataset is not as clean as ESC-50, particularly the ‘silence’ class where some samples contain the sound of bird calls, despite the existence of a bird class. Because of this, the dataset is considered harder to learn.
IV-A3 ARCA23K-FSD
ARCA23K-FSD [53] is a single-label dataset for environmental audio classification derived from the multi-label FSD50K environmental audio dataset, which in turn is collected from Freesound. The dataset contains 70 unbalanced classes and over 23,000 clips. FSD50K is manually human-labelled and contains clips varying in length from 0.3 seconds to 30 seconds. Due to the large number of classes, a similar approach to ESC-50 can be undertaken where k-fold cross-validation can be used for evaluation.
IV-A4 UrbanSound8k
UrbanSound8k [54] is an environmental sound classification dataset collected from Freesound. Each clip is no more than 4 seconds long, and a total of 8,732 clips are in the dataset. There are 10 unbalanced classes. The small amount of classes is problematic for zero-shot learning, as even splitting into three sets (train, validation and test) results in very little in each partition. An alternative is to use only train and test partitions.
IV-A5 TAU Urban Acoustic Scenes 2019
TAU 2019 [55] is an acoustic scene dataset where the aim isn’t to classify specific sounds but rather the type of location the sound is from, for example, a park or an airport. Each audio segment is 10 seconds long and belongs to one of 10 acoustic scene classes. Each scene has 1,440 audio clips. The audio clips in the dataset were collected by the authors rather than scraped or curated from existing sources online. This positions it as a useful dataset due to its unique audio clips, whereas the Freesound datasets have overlap due to their shared source. Like UrbanSound8k, there are only 10 classes so splitting the data into two sets (train and test) is more reasonable than adding a validation partition.
IV-A6 GTZAN
GTZAN [56] is a music genre dataset of 10 genres with 100 audio files for each class. Each clip is 30 seconds long. While this is an adjacent domain to environmental sound classification, it provides diversity to the experiments when many of the other datasets come from the same source (Freesound). The number of classes is equal to TAU 2019 and UrbanSound8k, motivating the use of the same partitioning strategy.
Partition Classes
ESC-50 fold 0 2, 3, 27, 29, 31, 35, 38, 40, 46, 48
fold 1 13, 19, 21, 22, 26, 32, 36, 39, 42, 49
fold 2 4, 10, 14, 17, 23, 24, 30, 33, 41, 45
fold 3 1, 6, 7, 18, 20, 25, 28, 34, 44, 47
fold 4 0, 5, 8, 9, 11, 12, 15, 16, 37, 43
ARCA23K-FSD fold 0 crash cymbal, run, zipper (clothing), acoustic guitar, gong, knock, train, crack, cough, cricket
fold 1 electric guitar, chewing mastication, keys jangling, female speech woman speaking, crumpling crinkling, skateboard, computer keyboard, bass guitar, stream, toilet flush
fold 2 tap, water tap faucet, squeak, snare drum, finger snapping, walk footsteps, meow, rattle (instrument), bowed string instrument, sawing
fold 3 rattle, slam, whoosh swoosh swish, hammer, fart, harp, coin (dropping), printer, boom, giggle
fold 4 clapping, crushing, livestock farm animals working animals, scissors, writing, wind, crackle, tearing, piano, microwave oven
fold 5 trumpet, wind instrument woodwind instrument, child speech kid speaking, drill, thump thud, drawer open close, male speech man speaking, gunshot gunfire, burping eructation, splash splatter
fold 6 female singing, wind chime, dishes pots pans, scratching (performance technique), crying sobbing, waves surf, screaming, bark, camera, organ
FSC22 train 0, 1, 2, 3, 4, 10, 11, 14, 16, 19, 20, 24, 25
val 6, 8, 9, 12, 13, 18, 22
test 5, 7, 15, 17, 21, 23, 26
UrbanSound8K train 0, 1, 2, 4, 5, 7, 8
val 3, 6, 9
TAU 2019 train 2, 3, 4, 5, 7, 8, 9
val 0, 1, 6
GTZAN train 0, 1, 2, 6, 7, 8, 9
val 3, 4, 5
TABLE I: Classes in each partition. For ESC-50, random partitions are used from previous work for consistency [19]. All other datasets have not been benchmarked for zero-shot learning before, so new partitions are created using a random number generator.
IV-B Embeddings
A modified YAMNet model [20] - a YAMNet architecture with layers removed and an embedding layer added in before the output layer - is trained using only the seen classes in each dataset partition. For all datasets, the embedding layer size is 128, the learning rate is $10^{-3}$, and the batch size is 16. The models are trained with early stopping and a patience of 200. This model is then used to retrieve audio embeddings to train each zero-shot learning method.
The word embedding model used is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/. Each method uses the same embedding vectors for each respective class and data sample.
IV-C Non-linear ALE
The non-linear ALE compatibility network has been used to great success in environmental audio [48, 20]. It represents the state-of-the-art method in zero-shot learning for environmental audio on ESC-50. Therefore, it is important to include it in this comparison.
The non-linear compatibility method has few hyperparameters and the parameters used by Xie et al. [48] for ESC-50 were repeated in our work for all datasets, except FSC22 which was tuned to train for 30 epochs.
IV-D CADA-VAE
Cross and distribution aligned variational autoencoder (CADA-VAE) [15] is a successful zero-shot learning method in computer vision. For this reason, it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
Hyperparameter tuning was conducted by starting with the default parameters from the computer vision experiments and manually tuning based on the validation sets. Both datasets used the same hyperparameters, aside from the number of classifier epochs. Overall, hyperparameter tuning was more useful in stabilising training and reducing the standard deviation than improving the accuracy. It was found that the large amount of hyperparameters made this method difficult and time-consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
The final training configuration was determined iteratively, guided by validation performance. The VAE is trained with warmup strategies for the Kullback–Leibler divergence (KLD), the cross-reconstruction error, and the distance error. The KLD has a factor of one with a warmup over epochs 0 to 93, the cross-reconstruction has a factor of 2.37 with a warmup over epochs 21 to 75, and the distance error has a factor of 5.13 with a warmup over epochs 6 to 22. The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs. The classifier uses a learning rate of $10^{-3}$ and a batch size of 32. It is trained for 23 epochs on ESC-50; 20 epochs on FSC22, UrbanSound8K and GTZAN; 30 epochs on TAU 2019; and 50 epochs on ARCA23K-FSD, with the number of epochs selected based on the size of each dataset. The loss of the VAE is a mean absolute error and the loss of the classifier is a negative log-likelihood loss. Both use an Adam optimiser. The size of the latent space is 64.
IV-E LisGAN
Leveraging invariant side generative adversarial network (LisGAN) [16] is a more recent GAN method that reports comparatively high results amongst GANs in zero-shot learning. LisGAN is a transductive method while we consider only the inductive case in these experiments. The transductive component is removed and the core idea of the method remains. A conditional Wasserstein GAN [38] is the main GAN method found in zero-shot learning. However, despite efforts to improve the method, it is still notoriously difficult to train [57, 43]. Smaller experiments were performed with a basic conditional GAN with no success, and often results were no better than random with a high variance across runs. LisGAN has shown to be significantly more stable and competitive amongst other zero-shot learning methods, even with the transductive component removed.
Hyperparameter tuning of LisGAN involved simple manual tuning, with only a small amount of parameters to tune. These were modified to suit the smaller audio datasets, as the datasets in computer vision are significantly larger. All datasets used the same hyperparameters.
LisGAN is used without retraining on unseen classes. The GAN is trained with a batch size of 64, 128 units in the hidden layer of the discriminator and generator, 50 epochs and a learning rate of $10^{-4}$. The classifier is trained with a learning rate of $10^{-2}$ and 50 epochs. The generator, discriminator and classifier use Adam optimisers. The classifier uses a negative log-likelihood loss.
ALE LisGAN CADA-VAE ZeroDiffusion
ESC-50 fold 0 $0.2278\pm 0.0091$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2605}}\pm 0.0227$ $\textbf{0.2915}\pm 0.0180$ $0.2598\pm 0.0140$
fold 1 $\textbf{0.4932}\pm 0.0093$ $0.4039\pm 0.0220$ $0.3980\pm 0.0181$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4665}}\pm 0.0194$
fold 2 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2937}}\pm 0.0129$ $0.2582\pm 0.0187$ $0.2500\pm 0.0251$ $\textbf{0.3110}\pm 0.0082$
fold 3 $0.2680\pm 0.0117$ $0.2150\pm 0.0195$ $\textbf{0.3013}\pm 0.0251$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2975}}\pm 0.0127$
test $0.2340\pm 0.0167$ $\textbf{0.2448}\pm 0.0165$ $0.2095\pm 0.0337$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2343}}\pm 0.0118$
average $0.3033\pm 0.0119$ $0.2765\pm 0.0199$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3102}}\pm 0.0240$ $\textbf{0.3138}\pm 0.0132$
ARCA23K-FSD fold 0 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2420}}\pm 0.0032$ $0.2183\pm 0.0128$ $0.2398\pm 0.0125$ $\textbf{0.2879}\pm 0.0153$
fold 1 $0.2876\pm 0.0028$ $0.3059\pm 0.0107$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3142}}\pm 0.0149$ $\textbf{0.3418}\pm 0.0150$
fold 2 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3723}}\pm 0.0061$ $0.3086\pm 0.0121$ $0.3164\pm 0.0120$ $\textbf{0.4819}\pm 0.0188$
fold 3 $\textbf{0.2546}\pm 0.0029$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2113}}\pm 0.0148$ $0.2036\pm 0.0101$ $0.1885\pm 0.0200$
fold 4 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2283}}\pm 0.0020$ $0.2173\pm 0.0113$ $0.2126\pm 0.0140$ $\textbf{0.2833}\pm 0.0279$
fold 5 $0.3205\pm 0.0050$ $0.3196\pm 0.0142$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3567}}\pm 0.0144$ $\textbf{0.4335}\pm 0.0321$
test $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2129}}\pm 0.0051$ $0.1772\pm 0.0068$ $\textbf{0.2414}\pm 0.0145$ $0.2081\pm 0.0261$
FSC22 val $0.2650\pm 0.0088$ $0.2990\pm 0.0219$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3179}}\pm 0.0137$ $\textbf{0.3269}\pm 0.0113$
test $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2968}}\pm 0.0070$ $0.2142\pm 0.0416$ $\textbf{0.3290}\pm 0.0339$ $0.2775\pm 0.0071$
UrbanSound8k $\textbf{0.4997}\pm 0.0071$ $0.3848\pm 0.0238$ $0.4741\pm 0.0808$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4961}}\pm 0.0777$
TAU2019 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4377}}\pm 0.004$ $0.3878\pm 0.0188$ $0.3158\pm 0.0248$ $\textbf{0.4857}\pm 0.0395$
GTZAN $0.5705\pm 0.0109$ $\textbf{0.5824}\pm 0.0287$ $0.5650\pm 0.0463$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.5803}}\pm 0.0281$
Average $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3227}}\pm 0.0076$ $0.2936\pm 0.0187$ $0.3137\pm 0.0242$ $\textbf{0.3486}\pm 0.0221$
TABLE II: Average accuracy and standard deviation over ten randomly seeded runs for the traditional non-linear ALE method and the three generative methods for the six chosen datasets. Our method, ZeroDiffusion, is similar or superior in almost all cases, with the exception of ESC-50 where no one method consistently performs the best. When averaging ESC-50 results, ZeroDiffusion is ahead of the other methods. The best accuracy in each row is bolded and the second-best accuracy is bolded in grey. The average value across all datasets is displayed in the final row, where ZeroDiffusion has the highest average accuracy and ALE reports the lowest overall variance.
IV-F ZeroDiffusion
ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in Section III-C. These experiments aim to determine the performance of this new method in the context of environmental audio. The existing success of diffusion in other areas of machine learning research motivates the creation and experimentation of this method.
The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 50 epochs. The number of samples generated for each unseen class corresponds to the average number of samples per class in the dataset. For TAU 2019, this is 1,440 samples per class; UrbanSound8k is unbalanced but averages 873 samples per class; GTZAN is 100 samples per class; ESC-50 uses 40 samples per class; and finally, ARCA23K-FSD is unbalanced and averages 339 per class. We train the non-linear compatibility network with a learning rate of $10^{-4}$, batch size of 64, and 10 epochs. For the diffusion loss terms, reconstruction loss was weighted $1.0$, maximum mean discrepancy at $1.0$, variance loss as $0.1$, mean reconstruction loss at $0.2$, and cosine similarity at $2.0$.
These hyperparameters were chosen using manual tuning across all validation sets. The same parameters were used for all experiments except for the synthetic dataset size, which changed relative to the dataset size.
V Results
Table II presents the results from experiments with the four methods - non-linear ALE, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50, ARCA23K-FSD, FSC22, UrbanSound8k, TAU 2019 and GTZAN.
ZeroDiffusion maintains an equally top result for UrbanSound8k and GTZAN, shared with ALE and LisGAN, while it outperforms other methods on TAU 2019. Notably, despite not outperforming all methods across all three smaller datasets, it is the only method to maintain a top result across all three. For the FSC22 dataset, the results are worse for ZeroDiffusion but it outperforms other methods on the validation set and underperforms on the test set. This indicates an issue with reliability for the new method; however, it is clear that the other presented methods also lack reliability.
Overall, the accuracy improvement of ZeroDiffusion across most datasets and partitions supports the use of diffusion models in the field of zero-shot learning. Adjacent works with Stable Diffusion [44, 45] are supported by ZeroDiffusion’s performance, indicating that diffusion methods more broadly have potential in a range of domains for zero-shot learning. ZeroDiffusion is agnostic to the specific data type and works entirely with embeddings, similarly to ALE, and therefore is a more flexible method than the image domain-specific Stable Diffusion methods.
V-A Comparison of Dataset Sizes
ARCA23K-FSD is a significantly larger dataset than any of the other datasets, and as such allows for a deeper 7-fold analysis. ZeroDiffusion outperforms the other methods in almost every case, and in the cases of folds 0, 2, 4 and 5 it has a significant improvement over any other method. This result indicates that ZeroDiffusion may draw benefits from the larger dataset, with richer embeddings produced by the diffusion model as a result of more classes available. However, despite the overall success, fold 3 extremely underperforms and could provide an avenue for improvement.
Whilst ESC-50 is not as large as ARCA23K-FSD, its large class size allows for 5-fold partitioning. However, the results for this dataset are extremely unstable when considering the best-performing methods. No one method consistently outperforms this dataset, and all methods outperform the others on at least one partition. To gain a better understanding, the average across all folds is calculated and ZeroDiffusion marginally has the highest accuracy. An indication of why this is happening may be revealed in the average number of audio clips per class in each dataset. ESC-50 has the lowest with 40 samples per class. This may directly result in the method failing to generalise. FSC22 is the next largest dataset per class with 75 samples, and this dataset similarly experienced issues with performance. Of the smaller datasets, TAU 2019 has 1,440 per class, the highest of any dataset in the experiments, and ZeroDiffusion produced a significant improvement over other methods. ARCA23K-FSD has on average 339 samples per class, a reasonable increase over both ESC-50 and FSC22, and reports good results. Given this is an unbalanced dataset, this may provide insight into why ZeroDiffusion underperforms on some folds in ARCA23K-FSD.
V-B Ablation
Notably, ZeroDiffusion incorporates the non-linear model in its methodology. The ALE results therefore represent an ablation of ZeroDiffusion without the diffusion-generated data component. The only other changes present are the weight decay term on the classifier ($10^{-4}$ for ALE and $10^{-5}$ for ZeroDiffusion) and the loss function (WARP for ALE and cross-entropy for ZeroDiffusion’s classifier); however, tests were done to align these with poorer results on both ALE and ZeroDiffusion across datasets when changing either or both terms. Given all of this, is it reasonable to conclude that the addition of a diffusion model to generate synthetic data for classifier training is responsible for the improved accuracy between ALE and ZeroDiffusion.
V-C Variance
Standard deviation is recorded for each experiment, and it is clear that across runs the generative-based methods have a high standard deviation. Meanwhile, ALE has a relatively low standard deviation. This indicates that generative-based methods, including ZeroDiffusion, are more variable and potentially unreliable compared to the more stable ALE method. Additionally, generative methods typically have a wide range of hyperparameters. Methods such as ALE have a low number of hyperparameters, making them user-friendly and simple to apply to new problems.
VI Discussion
Given the results of this four-model and six-dataset study, three areas are considered: model consistency and the importance of consistency in zero-shot learning; the impact of dataset size; and the issue of theoretical guarantees in zero-shot learning.
VI-A Consistency of Methods
In a true zero-shot learning real-world application, it is hard to know how well a model will perform on the target class due to a lack of available samples. Therefore, consistency across training runs and datasets is important for building trust in the model. The issue of the black-box nature of machine learning models hinders application to key global problems, and the added uncertainty of model performance further exasperates these problems.
When considering the standard deviation of each result in Table II, there is a high variance trend amongst the generative methods (LisGAN, CADA-VAE, ZeroDiffusion), with CADA-VAE reaching a high 8.08% standard deviation on the UrbanSound8k dataset and ZeroDiffusion not far behind with a 7.77% deviation. Meanwhile, the highest standard deviation for ALE is around 1% on ESC-50 partitions and GTZAN. This is a serious problem for some domains and indicates a need to develop more stable generative methods to make use of the improved accuracy of these methods.
Between datasets, there is also an inconsistency, as noted with ESC-50 having top results across partitions from every method used in the presented experiments. ZeroDiffusion has high accuracy on ARCA23K-FSD; however, it suffers issues with fold 3 and the test fold despite large improvements on the other folds. Prediction of model performance across partitions and datasets can better inform the application and improvement of these methods.
VI-A1 Data Diversity
From the results presented, larger datasets typically faired better with ZeroDiffusion. TAU 2019 may have only ten classes, but with 1,440 samples per class it represents a large audio dataset and this appears to result in an accuracy improvement over other methods. The smaller two datasets, ESC-50 (40 samples per class) and FSC22 (75 samples per class) experienced instability in performance across methods. ARCA23K-FSD provided a balance between the number of classes (70) and the number of samples per class (on average 339 samples), and the methods were more consistent in their performance, with ZeroDiffusion dominating. The dataset size, both in the number of classes and samples per class, may influence which zero-shot classification method is more suitable for individual scenarios and tasks. This is an important consideration when applying zero-shot learning in practice.
VI-A2 Theoretical Guarantees
Zero-shot learning lacks theoretical foundations, with methods developing through empirical results, as done in the presented work. There is little understanding of upper bounds on accuracy for a given set of unseen and seen classes. A given seen class set will not necessarily have enough information to extrapolate to particular unseen classes. A trivial example is if the seen class set is the empty set, then there is no information to learn. One indication of these theoretical guarantees is the differences between dataset sizes in the presented experiments and the effect on model stability. A deeper understanding of these requirements may improve overall classification performance.
VII Conclusion
This study presents a comparison of three generative methods and a baseline method in zero-shot environmental sound classification. Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly to the baseline and present higher variance.
The third generative method, ZeroDiffusion, is our novel diffusion-based method for zero-shot learning. It has a higher accuracy performance than the other three methods in most cases on the six chosen datasets, with a new state-of-the-art result on ARCA23K-FSD and TAU Urban Acoustics 2019, and reasonable results on ESC-50, FSC22, UrbanSound8k and GTZAN. The inclusion of TAU 2019 and GTZAN shows that the method can perform in applications other than environmental sound classification, with acoustic scene classification and music genre classification represented by these datasets. Future work on diffusion-based zero-shot models would benefit from considering other domains, such as image classification.
Acknowledgements
This research is supported by an Australian Government Research Training Program Scholarship to the first author.